OpenStudy (anonymous):

Another solving of equations. x + 3 x + 2 1 ----- = ----- + ---- x - 2 x - 2 2x

OpenStudy (anonymous):

x+3 x+2 1 --- - ---- = --- x-2 x-2 2x 1 1 -- = ------- x-2 2x 2x=x-2 x=-2

OpenStudy (anonymous):

\[2x(x+3) = 2x(x+2)+x-2\] \[2x ^{2}+6x= 2x ^{2}+4x+x-2\] \[x=-2\]